# Tightening the reins on non-minimal dark sector physics: Interacting Dark Energy with dynamical and non-dynamical equation of state

### v1

@@ I Introduction
On the other hand, from a more observational standpoint, IDE cosmology has emerged as a possible solution to cosmological tensions Salvatelli et al. (2014); Kumar and Nunes (2016); Xia and Wang (2016); Kumar and Nunes (2017a); Yang et al. (2017); Feng et al. (2019); Yang et al. (2019a, 2018b, 2018c); Li et al. (2019); Kumar et al. (2019); Pan et al. (2019a); Di Valentino et al. (2017); Yang et al. (2017); Feng et al. (2019); Yang et al. (2019a, 2018b, 2018c); Li et al. (2019); Kumar et al. (2019); Pan et al. (2019a); Yang et al. (2019b); Pan et al. (2019b); Benetti et al. (2019); Di Valentino et al. (2020a, b); Zhai et al. (2023); Pan and Yang (2023); Benisty et al. (2024). Allowing an exchange of energy-momentum from DM to DE, can increase the value of $H_{0}$ inferred from CMB observations, possibly restoring the agreement with the direct measurement provided by the SH0ES collaboration. In addition, as recently shown by some of us in Ref. Zhai et al. (2023), IDE seems to be supported by different independent CMB experiments, leading to an overall consistency of view concerning the amount of energy-momentum transferred in the dark sector. Having said that, whether or not this model could provide a successful solution for the Hubble trouble is still a matter of debate. The model suffers from the typical problem faced by any late-time solution (i.e., solutions that are aimed to solve the Hubble tension by introducing new physics post-recombination). Namely, Baryon Acoustic Oscillation (BAO) data and distance moduli measurements for Supernovae (SN) are very constraining on local distances, leaving us with little freedom to introduce any deviation from a basic $\Lambda$CDM late-time cosmology Keeley and Shafieloo (2023). As a result, when considering low-redshift ($z$) probes, the ability of IDE to increase the present-day expansion rate of the Universe is strongly reduced, if not completely lost. [^2] It is worth mentioning some caveats surrounding the use of BAO data. Firstly, volumetric BAO data might retain a residual model dependence from the template used in the analysis pipeline. In addition, as pointed out by some of us in Ref. Bernui et al. (2023), volumetric BAO data produce somewhat conflicting constraints on IDE compared to transverse 2D-BAO measurements, providing another element of concern.
@@
  • In reference to the previous point, a somewhat technical follow-up motivation involves noting that while dynamic behaviors have already been partially explored for different models (with old datasets), a proper treatment of $w(z)$ at different $z$ in the CPL parameterization is also missing. [^3] For instance, some of us in Ref. Yang et al. (2023) considered the so-called early-CPL (CPLe) parameterization in the context of similar yet different IDE cosmologies. The usual approach performed in previous studies on this topic consisted of imposing priors on the free parameters appearing in the CPL parameterization, potentially forcing the model deep into the phantom or quintessence regime at $z>0$. Instead, the peculiarity of our approach will be to ensure that $w(z)$ can always lie in the quintessence ($w(z)>-1$) or phantom ($w(z)<-1$) regime at any $z$, without the need to assume priors. This allows for proper sampling of the parameter space which has never been performed in the literature whatsoever. We will provide more details on this point in section III.
@@ II Interacting Dark Energy Cosmology
where $\xi$ represents the coupling parameter, and $\mathcal{H}$ denotes the conformal Hubble parameter, following the standard definition as in General Relativity: $3\mathcal{H}^{2}=8\pi G\;a^{2}(\eta)\sum\rho_{i}$, where $\rho_{i}$ represents the energy density of the $i$-th fluid. In addition to CDM and DE, we have also accounted for the presence of baryons, radiation, and neutrinos, including one massive and two massless species. Hence, consistent with the sign convention of $Q(t)$, $\xi>0$ ($<0$) denotes the transfer of energy from DE to CDM (from CDM to DE), respectively.
@@
Ensuring the stability of linear perturbations over time is crucial within the dynamical scenarios considered for the dark coupling in this study. As demonstrated in Gavela et al. (2009), the parameter known as the "doom factor," denoted as $\mathrm{d}=Q/(3\mathcal{H}(1+w_{x}))$, plays a pivotal role in determining the stability of the scalar modes. When $\mathrm{d}<0$, indicating stability, it implies that for $Q>0$, the equation of state parameter $w_{x}$ must be less than -1. Conversely, for $Q<0$, $w_{x}$ needs to be greater than -1 to maintain stability. In the following section, we will delve into a detailed description of the parameter space that ensures an absence of instabilities.
@@ III.1 Markov Chain Monte Carlo Analysis
  • Non-Dynamical DE EoS – In this case $w_{0}$ becomes an additional free parameter that we vary in two different regimes: the quintessence regime where $w_{0}>-1$ and the phantom regime where $w_{0}<-1$. To avoid instabilities in primordial perturbations, the coupling parameter $\xi$ can vary within the following priors: $\xi<0$ when $w_{0}>-1$ and $\xi>0$ when $w_{0}<-1$. Therefore the behavior of the DE EoS fixes the direction of energy-momentum transfer between DM and DE. In summary:
@@
    i) The quintessence regime where $w(z)>-1$ for any $z$. In the MCMC analysis, we sample over these two parameters, ensuring that for every randomly sampled pair of values $w_{0}$-$w_{a}$, the condition $w(z)>-1$ holds true at any $z$. If this condition is not met, the point is rejected. The test is performed dynamically during the MCMC run, without assuming any restrictive prior on the parameter $w_{a}$ controlling the dynamical evolution of $w(z)$. In contrast, we assume a prior $w_{0}>-1$ that automatically follows from requiring $w(z)>-1$ at $z=0$. Our methodology ensures proper sampling of the parameter space and convergence of the chains.
@@ IV.1.1 CMB-only
Considering only CMB temperature, polarization, and lensing spectra, we have limited power to simultaneously constrain the DE EoS $w_{0}$ and the coupling parameter $\xi$. Concerning the former, from P18, we obtain an upper limit of $w_{0}<-0.787$ at 95% CL. Replacing P18 with ACT data further relaxes this upper limit to $w_{0}<-0.56$. As for the coupling $\xi$, in both experiments we observe a 68% preference for non-vanishing energy-momentum flow ($\xi=-0.39^{+0.36}_{-0.13}$ from P18 and $\xi=-0.39^{+0.15}_{-0.30}$ from ACT). However, the preference vanishes at 95% CL. Overall, the two datasets lead to a lower bound on the coupling $\xi\gtrsim-0.8$, which is very broad and not particularly informative.
@@ IV.1.2 CMB and SN
As a next step, to gain some constraining power, we incorporate SN data into our analysis. When dealing with SN, a decision needs to be made regarding whether to consider the uncalibrated dataset or introducing the SH0ES calibration for the absolute SN magnitude. On the one hand, a conservative approach would involve using the uncalibrated Pantheon-Plus dataset and examining it alongside CMB measurements. On the other hand, if we take the results of the CMB-only analysis at face value, the Hubble tension is significantly reduced (mainly due to larger uncertainties). Therefore, using the SH0ES calibration is an alternative compelling decision. To illustrate the effects resulting from the supernova calibration, here we explore both of these possibilities.
@@
Having said that, it is worth considering that the situation looks very different when SN are calibrated with SH0ES. In this case, the constraints on the DE EoS become more restrictive on deviations away from the cosmological constant. We get $w_{0}<-0.843$ and $w_{0}<-0.775$ for P18+SN+SH0ES and ACT+SN+SH0ES, respectively. In addition, from both P18+SN+SH0ES ($\xi=-0.37^{+0.18}_{-0.11}$) and ACT+SN+SH0ES ($\xi=-0.49^{+0.13}_{-0.25}$), we observe a preference for a non-vanishing energy-momentum flow that is in remarkable agreement for the two experiments and persists at 95% CL. This preference for a non-vanishing interaction produces higher values of the Hubble parameter ($H_{0}=69.75\pm 0.67$ km/s/Mpc and $H_{0}=69.96\pm 0.70$ km/s/Mpc for Planck+SN+SH0ES and ACT+SN+SH0ES, respectively). In this case, the Hubble tension is reduced down to $\sim 2.5-2.7\sigma$.
In conclusion, combining CMB observations with uncalibrated SN fails to increase the expansion rate of the universe. However, the calibrated SN dataset could leave enough room to mitigate (even if not completely solving) the Hubble constant tension.
@@ IV.1.3 CMB and BAO
In conclusion, including BAO data, no room is left to solve the Hubble tension.
@@ V.1.1 CMB-only
First and foremost, we note that all the concerns we pointed out in the non-dynamical case about the geometrical degeneracy among cosmological parameters observed for the CMB-only case apply also in the dynamical case. This problem can become even more relevant since we now have one additional parameter, $w_{a}$, describing the redshift evolution of $w(z)$. With this premise, it is not surprising that for P18 we are not able to constrain $w_{0}$. However, we can get a 95% lower limit on $w_{a}>-1.22$. Similar results are obtained for ACT: $w_{0}$ is unconstrained while $w_{a}$ turns out to be $w_{a}>-1.21$ at 95% CL. Concerning the coupling parameter, we note that in the dynamical case, we get a preference for a non-vanishing interaction $\xi=-0.40^{+0.36}_{-0.15}$ for P18 and $\xi=-0.49^{+0.23}_{-0.28}$ for ACT. Also in the dynamical case, the two experiments agree about the possible amount of energy-momentum to be transferred from DM to DE. This preference towards an interacting dark sector produces higher values of the present-day expansion rate which reads $H_{0}=71.6\pm 2.5$ km/s/Mpc for P18 and $H_{0}=72.6^{+3.1}_{-2.7}$ km/s/Mpc for ACT. Therefore, despite having large uncertainties, the CMB-only case suggests that considering dynamical quintessence models can facilitate solving the Hubble tension compared to the respective non-dynamical case. However, to confirm this preference, it is mandatory to test the model against low-redshift data.
@@ V.2.2 CMB and SN
Combining CMB and SN together, we enhance the constraining power, narrowing down the upper limit on the coupling parameter $\xi$. Specifically, using uncalibrated SN, for P18+SN we obtain $\xi<0.224$, while for ACT+SN, we get $\xi<0.269$. Further tightening occurs when considering the SH0ES calibration. For P18+SN+SH0ES and ACT+SN+SH0ES, we find $\xi<0.0647$ and $\xi<0.103$, respectively. However, uncalibrated SN data yield lower values $H_{0}=66.31\pm 0.72$ km/s/Mpc for P18+SN and $H_{0}=66.14\pm 0.89$ km/s/Mpc for ACT+SN. These results are in clear tension with SH0ES. Even after calibrating SN with SH0ES, despite slightly higher values of the present-day expansion rate are obtained ($H_{0}=68.48\pm 0.48$ km/s/Mpc for P18+SN+SH0ES and $H_{0}=69.39\pm 0.63$ km/s/Mpc for ACT+SN+SH0ES), the tension persists.
@@ VI Conclusions
Overall, our comprehensive reanalysis shows that updated BAO and SN data appear to constrain the possibility that $w_{0}$ or $w(z)$IDE alone can conclusively resolve the Hubble tensionfor the proposed interaction model. Using the SH0ES calibration for SN (which is a well-motivated choice given the larger values of $H_{0}$ produced by the CMB-only analysis), we still have room to mitigate, although not fully solve, the Hubble trouble. Additionally, some scatter combinations of data involving BAO also lead to a higher present-day expansion rate for the phantom case. However, considering the joint CMB+SN+BAO measurement, we always settle down to values of $H_{0}$ similar to those inferred within a $\Lambda$CDM-like late-time cosmology. It is worth investigating whether some new ingredients could be added to the IDE cosmology, which can overcome the SN and BAO issues pointed out in this article.

### v2

@@ I Introduction
On the other hand, from a more observational standpoint, IDE cosmology has emerged as a possible solution to cosmological tensions Salvatelli et al. (2014); Kumar and Nunes (2016); Costa et al. (2017); Xia and Wang (2016); Kumar and Nunes (2017a); Yang et al. (2017); Feng et al. (2019); Yang et al. (2019a); Zhang et al. (2018); Yang et al. (2018b, c); Li et al. (2019); Kumar et al. (2019); Pan et al. (2019a); Di Valentino et al. (2017); Yang et al. (2017); Feng et al. (2019); Yang et al. (2019a, 2018b, 2018c); Li et al. (2019); Kumar et al. (2019); Pan et al. (2019a); Yang et al. (2019b); Pan et al. (2019b); Benetti et al. (2019); Di Valentino et al. (2020a, b); Liu et al. (2022); Zhao et al. (2023); Zhai et al. (2023); Pan and Yang (2023); Benisty et al. (2024); Giarè et al. (2024); Sabogal et al. (2024). Allowing an exchange of energy-momentum from DM to DE, can increase the value of $H_{0}$ inferred from CMB observations, possibly restoring the agreement with the direct measurement provided by the SH0ES collaboration. In addition, as recently shown by some of us in Ref. Zhai et al. (2023), IDE seems to be supported by different independent CMB experiments, leading to an overall consistency of view concerning the amount of energy-momentum transferred in the dark sector. Having said that, whether or not this model could provide a successful solution for the Hubble trouble is still a matter of debate. The model suffers from the typical problem faced by any late-time solution (i.e., solutions that are aimed to solve the Hubble tension by introducing new physics post-recombination). Namely, Baryon Acoustic Oscillation (BAO) data and distance moduli measurements for Supernovae (SN) are very constraining on local distances, leaving us with little freedom to introduce any deviation from a basic $\Lambda$CDM late-time cosmology Krishnan et al. (2021); Keeley and Shafieloo (2023). As a result, when considering low-redshift ($z$) probes, the ability of IDE to increase the present-day expansion rate of the Universe is strongly reduced, if not completely lost. [^2] It is worth mentioning some caveats surrounding the use of BAO data. Firstly, volumetric BAO data might retain a residual model dependence from the template used in the analysis pipeline. In addition, as pointed out by some of us in Ref. Bernui et al. (2023), volumetric BAO data produce somewhat conflicting constraints on IDE compared to transverse 2D-BAO measurements, providing another element of concern.
@@ II Interacting Dark Energy Cosmology
where $\xi$ represents the dimensionless coupling constant which is independent of cosmic time. Here, $\mathcal{H}$ denotes the conformal Hubble parameter, following the standard definition as in General Relativity: $3\mathcal{H}^{2}=8\pi G\;a^{2}(\eta)\sum\rho_{i}$, where $\rho_{i}$ represents the energy density of the $i$-th fluid. In addition to CDM and DE, we have also accounted for the presence of baryons, radiation, and neutrinos, including one massive and two massless species. Hence, consistent with the sign convention of $Q(t)$, $\xi>0$ ($<0$) denotes the transfer of energy from DE to CDM (from CDM to DE), respectively.  [^3] It is worth noting that Ref. Pan et al. (2020c) explicitly demonstrated how interaction rates, as considered in Eq. (3), can naturally arise from first principles when exploring well-motivated field theories for scenarios of IDE.
@@
Ensuring the stability of linear perturbations over time is crucial within the dynamical scenarios considered for the dark coupling in this study. As demonstrated in Gavela et al. (2009), the parameter known as the "doom factor", denoted as $\mathrm{d}=Q/(3\mathcal{H}(1+w_{x}))$, plays a pivotal role in determining the stability of the scalar modes. When $\mathrm{d}<0$, indicating stability, it implies that for $Q>0$, the equation of state parameter $w_{x}$ must be less than $-1$. Conversely, for $Q<0$, $w_{x}$ needs to be greater than $-1$ to maintain stability. In the following section, we will delve into a detailed description of the parameter space that ensures an absence of instabilities.
@@ III.1 Markov Chain Monte Carlo Analysis
[…]
@@
    i) The quintessence regime where $w(z)>-1$ for any $z$. In the MCMC analysis, we sample over these two parameters, ensuring that for every randomly sampled pair of values $w_{0}$-$w_{a}$, the condition $w(z)>-1$ holds true at any $z$ [^5] Previous studies typically imposed priors on the CPL parameterization’s free parameters to force the model into the phantom or quintessence regime; see, e.g., Ref. Yang et al. (2023), where some of us considered the CPL parameterization in similar yet distinct IDE cosmologies. In contrast, our approach ensures that $w(z)$ can always lie in the quintessence ($w(z)>-1$) or phantom ($w(z)<-1$) regime at any $z$ without assuming priors but by dynamically checking $w(z)$ during the MCMC. Although this may seem like a technical detail within the CPL parameterization (where both approaches yield similar results), it becomes crucial when studying dynamical dark energy models beyond CPL (which is not done in this study). In such cases, conditions on $w(z)$ cannot always be easily mapped into priors on parameters, and our method allows for proper sampling of the parameter space, which has not been explored in the literature. If this condition is not met, the point is rejected. The test is performed dynamically during the MCMC run, without assuming any restrictive prior on the parameter $w_{a}$ controlling the dynamical evolution of $w(z)$. In contrast, we assume a prior $w_{0}>-1$ that automatically follows from requiring $w(z)>-1$ at $z=0$. Our methodology ensures proper sampling of the parameter space and convergence of the chains.
@@ IV.1.1 CMB-only
Considering only CMB temperature, polarization, and lensing spectra, we have limited power to simultaneously constrain the DE EoS $w_{0}$ and the coupling parameter $\xi$. Concerning the former, from P18, we obtain an upper limit of $w_{0}<-0.787$ at 95% CL. Replacing P18 with ACT data further relaxes this upper limit to $w_{0}<-0.56$. As for the coupling $\xi$, in both experiments we observe a 68% CL preference for non-vanishing energy-momentum flow ($\xi=-0.39^{+0.36}_{-0.13}$ from P18 and $\xi=-0.39^{+0.15}_{-0.30}$ from ACT). However, this preference diminishes at the 95% confidence level, and the analysis of both datasets indicates a broad lower bound of $\xi\gtrsim-0.8$, which lacks specificity and informative value. As expected, the main challenge arises from the so-called geometrical degeneracy among parameters. Essentially, different combinations of late-time cosmic parameters can be adjusted to keep the acoustic angular scale $\theta_{s}$ – defined by the ratio of the comoving sound horizon at recombination to the comoving distance to last scattering – constant as long as both quantities change proportionally. As a result, measurements based solely on CMB data, which can accurately determine this scale, cannot impose strong constraints on the (dynamical) IDE model by themselves unless late-time data are also incorporated to partially break this degeneracy.
@@ IV.1.2 CMB and SN
As a next step, to gain some constraining power, we incorporate SN data into our analysis. When dealing with SN, a decision needs to be made regarding whether to consider the uncalibrated dataset or introducing the SH0ES calibration for the absolute SN magnitude. On the one hand, a conservative approach would involve using the uncalibrated Pantheon-Plus dataset and examining it alongside CMB measurements. On the other hand, if we take the results of the CMB-only analysis at face value, the Hubble tension is significantly reduced (mainly due to larger uncertainties). Therefore, using the SH0ES calibration is an alternative compelling decision in some specific cases. [^6] In this regard, we would like to clarify some important aspects concerning the tension among datasets. As already pointed out in the text, when CMB data are analyzed alone, $H_{0}$ is poorly constrained. Due to the large uncertainties, the value of $H_{0}$ inferred from CMB data in this model is not necessarily in disagreement with SH0ES. This allows us to legitimately calibrate SN with SH0ES and analyze P18+SN+SH0ES to see whether the SH0ES calibration is supported by the model (i.e., to what extent we can increase $H_{0}$). On the other hand, the Hubble tension can be reframed as a tension among calibrators: SN calibrated with SH0ES and BAO calibrated with CMB, assuming standard early-time (i.e., pre-recombination) cosmology (as in the IDE model), are in strong tension (see Fig. 1 of Ref Pogosian et al. (2022)). This means that combining CMB+BAO+SN+SH0ES would be inappropriate due to this tension. Indeed, we never consider such a combination of data. Conversely, when CMB+BAO+SN are analyzed together, the SH0ES calibration is never used. Therefore, we ensure that we never combine datasets that are in tension with each other while exploring all possible informative combinations of data.
@@
Having said that, it is worth considering that the situation looks very different when SN are calibrated with SH0ES. [^7] We would like to cautiously remark that CMB (P18/ACT)+SN and SN+SH0ES are in tension at more than $3\sigma$, highlighting the significant impact of assuming or not assuming a SH0ES calibration when dealing with SN measurements. In this case, the constraints on the DE EoS become more restrictive on deviations away from the cosmological constant. We get $w_{0}<-0.843$ and $w_{0}<-0.775$ for P18+SN+SH0ES and ACT+SN+SH0ES, respectively. In addition, from both P18+SN+SH0ES ($\xi=-0.37^{+0.18}_{-0.11}$) and ACT+SN+SH0ES ($\xi=-0.49^{+0.13}_{-0.25}$), we observe a preference for a non-vanishing energy-momentum flow that is in remarkable agreement for the two experiments and persists at 95% CL. This preference for a non-vanishing interaction produces higher values of the Hubble parameter ($H_{0}=69.75\pm 0.67$ km/s/Mpc and $H_{0}=69.96\pm 0.70$ km/s/Mpc for Planck+SN+SH0ES and ACT+SN+SH0ES, respectively).
In conclusion, combining CMB observations with uncalibrated supernovae does not lead to an increase in the measured expansion rate of the universe. However, using a calibrated supernova dataset may result in a slight increase in the Hubble constant which is primarily driven by the SH0ES calibration.
@@ IV.1.3 CMB and BAO
When we analyze SN and BAO separately, the impact of different geometric measurements on our primary parameters of interest becomes apparent. Specifically, the BAO sample plays a crucial role in breaking the statistical degeneracy within our extensive parameter space. In conclusion, including BAO data, no room is left to solve the Hubble tension and the coupling parameter is very well limited.
@@ IV.2.4 Joint Analyses
We conclude this section with an important final remark: the cases $w_{0}>-1$ and $w_{0}<-1$ do not necessarily produce the same magnitudes of the coupling parameter $\xi$, nor the same value of $H_{0}$ in the limit $\xi\to 0$. These discrepancies arise because the sign of $\xi$ induces different corrections in $H_{0}$ and $\Omega_{m}$, which combine with the well-known correlation between $w_{0}$ and $H_{0}$, differing in the quintessence and phantom regimes. As a result, due to the varying correlations introduced by $\xi$, the final outcomes of analyses with $w_{0}>-1$ and $w_{0}<-1$ should not be expected to match. However, from a physical point of view, this is not a cause for concern because the two regimes differ significantly in their physical nature and potential microphysical realizations. The sign of $\xi$ predicts distinct cosmological behaviors, featuring an energy-momentum flow in opposite directions, and this could fundamentally alter the model’s physical and theoretical characteristics.
@@ V.1.1 CMB-only
First and foremost, we note that all the concerns we pointed out in the non-dynamical case about the geometrical degeneracy among cosmological parameters observed for the CMB-only case apply also in the dynamical case. This problem can become even more relevant since we now have one additional parameter, $w_{a}$, describing the redshift evolution of $w(z)$. With this premise, it is not surprising that for P18 we are not able to constrain $w_{0}$. [^8] Given that the constraints on the DE EoS often reach the prior bounds, a few important remarks on the prior range adopted for $w_{0}$ and $w_{a}$ and their possible implications for constraints on relevant parameters such as $H_{0}$ and $\xi$ are in order. Firstly, we note that no significant correlation is found between the parameters describing the EoS, $H_{0}$, and $\xi$. As seen from Figure 3 and  Figure 4, the probability contours at 68% and 95% are essentially flat and do not show significant shifts in $\xi$ and $H_{0}$ when $w_{0}$ and $w_{a}$ are varied. Therefore, the choice of prior on the DE EoS is not crucial for the constraints on key parameters related to cosmological tensions. We explicitly tested this aspect by adopting different case study priors on $w_{0}$ and $w_{a}$ without observing any relevant changes in the results for $H_{0}$ and $\xi$ (in both quintessence and phantom cases). Notice also that, for direct comparison of the results, we are using the standard priors on all parameters that are widely adopted in the literature. However, we can get a 95% lower limit on $w_{a}>-1.22$. Similar results are obtained for ACT: $w_{0}$ is unconstrained while $w_{a}$ turns out to be $w_{a}>-1.21$ at 95% CL. Concerning the coupling parameter, we note that in the dynamical case, we get a preference for a non-vanishing interaction $\xi=-0.40^{+0.36}_{-0.15}$ for P18 and $\xi=-0.49^{+0.23}_{-0.28}$ for ACT. Also in the dynamical case, the two experiments agree about the possible amount of energy-momentum to be transferred from DM to DE. This preference towards an interacting dark sector produces higher values of the present-day expansion rate which reads $H_{0}=71.6\pm 2.5$ km/s/Mpc for P18 and $H_{0}=72.6^{+3.1}_{-2.7}$ km/s/Mpc for ACT. Therefore, despite having large uncertainties, the CMB-only case suggests that considering dynamical quintessence models can facilitate solving the Hubble tension compared to the respective non-dynamical case. However, to confirm this preference, it is mandatory to test the model against low-redshift data.
@@ V.2.2 CMB and SN
Combining CMB and SN together, we enhance the constraining power, narrowing down the upper limit on the coupling parameter $\xi$. Specifically, using uncalibrated SN, for P18+SN we obtain $\xi<0.224$, while for ACT+SN, we get $\xi<0.269$. Further tightening occurs when considering the SH0ES calibration. For P18+SN+SH0ES and ACT+SN+SH0ES, we find $\xi<0.0647$ and $\xi<0.103$, respectively. However, uncalibrated SN data yield lower values $H_{0}=66.31\pm 0.72$ km/s/Mpc for P18+SN and $H_{0}=66.14\pm 0.89$ km/s/Mpc for ACT+SN. These results are in clear tension with SH0ES.
@@ VI Conclusions
Overall, our comprehensive reanalysis shows that updated BAO and SN data appear to constrain the possibility that $w_{0}$ or $w(z)$IDE alone can conclusively resolve the Hubble tension for the proposed interaction model. Using the SH0ES calibration for SN (which is a well-motivated choice given the larger values of $H_{0}$ produced by the CMB-only analysis), we still have room to mitigate, although not fully solve, the Hubble trouble. Additionally, some scatter combinations of data involving BAO also lead to a higher present-day expansion rate for the phantom case. However, considering the joint CMB+SN+BAO measurement, we always settle down to values of $H_{0}$ similar to those inferred within a $\Lambda$CDM-like late-time cosmology. Yet another interesting aspect of our updated analysis is that the state-of-the-art constraints on the DE EoS are relaxed in models that feature a dynamic evolution of $w(z)$ when interactions in the dark sector of the cosmological model are considered. This is due to the increased dimensionality of the parameter space that introduces additional degeneracies among the parameters, leading to larger uncertainties. Therefore, a potential direction for future work could involve extending the analysis to additional observational probes, particularly those related to perturbations and the growth of structures (e.g., weak lensing), which are not accounted for in this work due to the complexity of treating nonlinear scales. Overall, it is worth investigating whether some new ingredients could be added to the IDE cosmology, which can overcome the SN and BAO issues pointed out in this article.